OpenStudy (anonymous):

-8/5 - 1/2y= -8/3 Can someone walk me through this? If I recall I need to change each number to a common denominator, just double checking I'm doing this right.

OpenStudy (faiqraees):

yes

OpenStudy (faiqraees):

but first you have to take the numbers on the right hand side of the equation and the variables on the left hand side.

OpenStudy (anonymous):

so my problem should start out as -8/5 -1/2 +8/3=y and then it should be -48/30 -15/30 +80/30=y correct? @FaiqRaees

OpenStudy (faiqraees):

no no see -1/2 is being multiplied with y you cant take that on r.h.s it should be like\[-1/2 *y=-8/3+8/5\]

OpenStudy (anonymous):

ahh ok... so after that step my problem should look like -15/30*y=-80/30+48/30 @FaiqRaees

OpenStudy (faiqraees):

yes

OpenStudy (anonymous):

so if i add them together I'm left with -15/30y=32/30... now do i just add the 32 to the 15 and simplify??? @FaiqRaees

OpenStudy (anonymous):

sorry id subtract @FaiqRaees

OpenStudy (faiqraees):

first of all there should be a negative side with 32 correct that first

OpenStudy (anonymous):

-15/30y=-32/30

OpenStudy (faiqraees):

yep now multiply 30 on both sides

OpenStudy (anonymous):

-450/30y=-960/30 @FaiqRaees

OpenStudy (anonymous):

so then I'm left with -15y=-32

OpenStudy (anonymous):

and final answer would be -32/15 correct?

OpenStudy (faiqraees):

multiply both sides with -1

OpenStudy (faiqraees):

15y = 32 y = 32/15

OpenStudy (anonymous):

so y=32/15

OpenStudy (faiqraees):

yes

OpenStudy (anonymous):

perfect. thank you! @FaiqRaees

OpenStudy (faiqraees):

No problem